A baseball player bats the ball at 60° with the ground at a velocity of 50m/s-1. A fielder immediately runs in the opposite direction at a speed of 7m/s-1 in time to catch the ball.


Calculate the distance travelled by the fielder.


1
Expert's answer
2021-04-05T11:12:27-0400

"t=\frac{2v sin \alpha}{g},"

"l=Vt=\frac{2Vv sin \alpha}{g},"

"l=\frac{2\cdot 7\cdot 50\cdot 0.87}{9.8}=62~m."